# Learning “Partner-Aware” Collaborators in Multi-Party Collaboration

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly bring deployed in agentic settings where they act as collaborators with humans. Therefore, it is increasingly important to be able to evaluate their abilities to collaborate effectively in multi-turn, multi-party tasks. In this paper, we build on the AI alignment and “safe interruptability” literature to offer novel theoretical insights on collaborative behavior between LLM-driven collaborator agents and an intervention agent. Our goal is to learn an ideal “partner-aware” collaborator that increases the group’s common-ground (CG)—alignment on task-relevant propositions—by intelligently collecting information provided in interventions by a partner agent.We show how LLM agents trained using standard RLHF and related approaches are naturally inclined to ignore possibly well-meaning interventions, which makes increasing group common ground non-trivial in this setting. We employ a two-player Modified-Action MDP to examine this suboptimal behavior of standard AI agents, and propose Interruptible Collaborative Roleplayer (ICR)—a novel “partner-aware” learning algorithm to train CG-optimal collaborators. Experiments on multiple collaborative task environments show that ICR, on average, is more capable of promoting successful CG convergence and exploring more diverse solutions in such tasks.
@@ 1 Introduction
We hypothesize that optimizing for general task utility (e.g., interventions that ultimately lead to correct task solutions) through counterfactual regularization encourages “partner-aware” behavior, leading to higher common ground convergence. Importantly, under our hypothesis, a true collaborator agent itself never has any more information than the aggregate of the group, and so common ground convergence should occur even without explicitly training for it. That is, an intentional collaborator learns to adapt: integrating helpful interventions while critically evaluating flawed ones. This ability to distinguish signal from noise fosters belief alignment as an emergent property of training, with practical benefits. In zero-shot or real-world collaborative settings, where intervention styles or partners are unfamiliar, counterfactually-trained agents should generalize better by leveraging learned notions of intervention quality. We validate this through a method we call Interruptible Collaborative Roleplayer (ICR), where we withhold common ground-based rewards during training and showing that such agents still achieve greater convergence than sophisticated LLM-agent training baselines, suggesting they have internalized collaboration principles transferable across partners and task to “in-the-wild” settings. Our work advances the state of the art in LLM-based collaborative agents through the following contributions [^1] Our code is available at https://github.com/csu-signal/ICR:
@@
  • Theoretical insights demonstrating why standard reinforcement learning and preference alignment algorithms (e.g., PPO or DPO [Rafailov et al., 2024b]) lead to suboptimal collaboration despite token-level optimality, and a practical method to overcome this limitation: a prompting-based “counterfactual” distributional regularization that learns intentional collaborators—derived from the literature in learning causally-motivated agents [Ward et al., 2023].
@@ 2 Related Work
While interruptibility has been studied in safety-critical RL [Orseau and Armstrong, 2016, Hadfield-Menell et al., 2017], it is equally vital in collaborative dialogue, where agents must discern whether interventions aid or hinder shared understanding [Grice, 1975, Sutton and Rao, 2024]. Prior work has explored these ideas in adversarial or game-theoretic contexts [Langlois and Everitt, 2021, Ward et al., 2023], but less so in multi-party deliberative language settings [Nath et al., 2025a, Obiso et al., 2025]. We extend this by training collaborator agents that are counterfactually robust—they update their beliefs when interventions are helpful, while resisting misleading or misaligned input.
@@ 3 The Collaborator’s Dilemma
An AI collaborator that merely mimics behavior patterns or reflexively adopts suggestions may initially appear cooperative but will demonstrate poor robustness when faced with interventions that are noisy, irrelevant, or potentially misleading [Jaques et al., 2019]. Rather, it needs to develop what Ward et al. [2023] terms “intentionality”—the capacity to autonomously evaluate interventions based on their causal impact on task outcomes rather than superficial plausibility. To address this limitation, we need a learning paradigm that enables collaborators to be partner-aware—capable of adapting to specific intervention agents through selective incorporation of helpful suggestions while maintaining invariance to misleading ones—thereby developing the "intentionality" necessary for robust collaborative reasoning. Such a collaborator would maintain reasoned agency in the face of various intervention qualities, leading to more robust collaboration and better common ground convergence across diverse interaction scenarios. In other words, effective collaborators must remain safely interruptible [Orseau and Armstrong, 2016]a delicate balance between receptive and robust that renders them open to incorporating valuable insights that genuinely contribute to task success, yet capable of maintaining their reasoning integrity when faced with misleading suggestions. This motivates our Interruptible Collaborative Roleplayer (ICR) learning algorithm.
@@ 4 Method: Interruptible Collaborative Roleplayer
Our counterfactual invariance approach directly addresses the suboptimality gap identified in Theorem 3.2. Initially during training of a collaborator policy $\pi_{C}^{CI}$ with counterfactual invariance regularization, the counterfactual KL divergence $\Delta_{\text{CF}}(\pi_{C}^{CI})=\mathbb{E}_{s,a^{I}}[D_{KL}(\pi_{C}^{CI}(\cdot|s,a^{I})\,\|\,\pi_{C}^{CI}(\cdot|s^{\text{CF}},a^{I}))]$ will be high as the policy has not yet learned to distinguish intervention quality, but decreases as training progresses and the policy acquires counterfactual robustness. As established in Theorem B.4, this directly bounds the suboptimality gap: $J_{\mathcal{M}}(\pi_{C}^{*})-J_{\mathcal{M}}(\pi_{C}^{CI})\leq\frac{2\gamma R_{max}}{(1-\gamma)^{2}}(\epsilon_{task}+C\cdot\Delta_{\text{CF}}(\pi_{C}^{CI}))$. Theoretically, as $\lambda_{\text{Intent}}\rightarrow\infty$, $\Delta_{\text{CF}}(\pi_{C}^{CI})$ approaches zero, making our policy’s performance approach that of the optimal policy $\pi_{C}^{*}$ (subject to task optimization constraints $\epsilon_{task}$). This theoretical guarantee connects directly to Lemma A.2 showing that while preference-aligned policies achieve Bellman optimality at both token and intervention levels in the underlying MDP, they remain suboptimal in the MAMDP due to failing to account for intervention quality. Our counterfactual invariance objective $\mathcal{J^{*}}(\theta_{C})$ bridges this gap by explicitly teaching collaborator LLM agents to distinguish between interventions based on their causal impact on task outcomes during training, rather than merely using their in-context learning capacity. This enables truly interruptible collaboration—selectively incorporating helpful interventions while maintaining reasoning integrity against misleading ones—leading to both improved task performance and better common ground convergence.
@@ 5 Experimental Design
For training data, we first collect MAMDP interaction trajectories (as defined in Sec. 3) on these two domains over 15 turns [^6] This roughly reflects the true distribution of back and forth interactions in the original DeliData task. using a high-capacity LLM (GPT-4o [OpenAI et al., 2024]) to roleplay both the intervener and the collaborator agents in each task. See Figs. 2 and 5 for prompts. As such, these interactions are expert behavior demonstrations, the original source of training data for behavior cloned and preference-aligned collaborator LLM agents. For evaluation, all trained ICR collaborators and competing baselines are first deployed following the MAMDP interaction in the expert data collection, and then evaluated, primarily on their ability to reach consensus during the collaboration. In all cases, we use a fixed intervention agent—an instance of GPT-4o prompted with the same system prompt in all evaluation runs—with $T=0$ and top-$p$ of 0.9 for sampling. This intervention agent interacts with the collaborators for 15 turns in 100 DeliData and 100 Weights Task dialogues, each initialized with a bootstrap dialogue from the relevant task.
@@
The “no-press” setting explores whether collaborator agents can achieve objective alignment on accurate decisions without explicit modeling of how interventions influence common ground formation. Therefore, to control for language interpretability, rather than using full natural language, collaborator agents act over a discrete space of structured beliefs—allowing us to evaluate grounded reasoning without requiring fluency. In the Weights Task, collaborators express beliefs as symbolic propositions over block weights (i.e., green > red, blue = 10g), while in DeliData, agents select from predefined stances—support, oppose, unsure, or consider_latertoward questions of which cards to flip. Agents are trained independently using only task-specific proxy rewards: factual accuracy in the Weights domain, and logically aligned card-checking in DeliData (e.g., +1 for supporting parts of a correct solution—a vowel or odd number, -1 for incorrect support, +0.5 for justified uncertainty). Using a proxy reward during training is intuitive as well as fair for baseline comparisons, since otherwise RL-based agent training is prone to reward hacking [^7] In fact, in our preliminary experimentation we found that rewarding agents with a consensus signal is counterproductive and often leads to reduced task-specific utility or correctness over propositions. [Strathern, 1997, Amodei et al., 2016]. In the no-press condition, evaluation follows an exact reward function $R(a^{c})$ that can be directly computed from the discrete solutions chosen by collaborators. For the full-press evaluation, we use an LLM-Judge [Zheng et al., 2023, Lambert et al., 2025]-based reward $R(s,a^{c})$, where $s$ is some dialogue context with the intervention present and $a^{c}$ are discrete actions inferred by the LLM-Judge based on the collaborator utterances in context.
@@ 6 Results and Analysis
We report “full-press” and “no-press” results for both tasks. For Weights Task (WTD), we report accuracy scores (ACC), a composite metric that multiplies the percentage of correct propositions by the total size of the common ground, rewarding both factual accuracy and common-ground convergence, while penalizing trivial solutions. For example, ACC of 14 indicates that the collaborator agents were able to recover 14 out a 37 total theoretically possible propositions at the end of the collaboration, adjusted for correctness. For DeliData, we report both accuracy (ACC)—a task-specific fine-grained score [Karadzhov et al., 2024] based on the final submission (after $N=15$ turns)—and common ground gain (CG), defined as the net increase in unique solution types introduced during the dialogue beyond those initially proposed. Specifically, we subtract the number of unique solution frameworks (e.g., $Odd$, VowelVowel, $Odd+Vowel$ in DeliData) initially proposed by the collaborator agents from the total number of distinct solutions considered throughout the entire dialogue. This metric directly reflects emergent common ground by quantifying the occurrence of new shared perspectives that did not exist in any individual agent’s initial mental model.

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly being deployed in agentic settings where they act as collaborators with humans. Therefore, it is increasingly important to be able to evaluate their abilities to collaborate effectively in multi-turn, multi-party tasks. In this paper, we build on the AI alignment and “safe interruptability” literature to offer novel theoretical insights on collaborative behavior between LLM-driven collaborator agents and an intervention agent. Our goal is to learn an ideal “partner-aware” collaborator that increases the group’s common-ground (CG)—alignment on task-relevant propositions—by intelligently collecting information provided in interventions by a partner agent. We show how LLM agents trained using standard RLHF and related approaches are naturally inclined to ignore possibly well-meaning interventions, which makes increasing group common ground non-trivial in this setting. We employ a two-player Modified-Action MDP to examine this suboptimal behavior of standard AI agents, and propose Interruptible Collaborative Roleplayer (ICR)—a novel “partner-aware” learning algorithm to train CG-optimal collaborators. Experiments on multiple collaborative task environments show that ICR, on average, is more capable of promoting successful CG convergence and exploring more diverse solutions in such tasks.
@@ 1 Introduction
We hypothesize that optimizing for general task utility (e.g., interventions that ultimately lead to correct task solutions) through counterfactual regularization encourages “partner-aware” behavior, leading to higher common ground convergence. Importantly, under our hypothesis, a true collaborator agent itself never has any more information than the aggregate of the group, and so common ground convergence should occur even without explicitly training for it. That is, an intentional collaborator learns to adapt: integrating helpful interventions while critically evaluating flawed ones. This ability to distinguish signal from noise fosters belief alignment as an emergent property of training, with practical benefits. In zero-shot or real-world collaborative settings, where intervention styles or partners are unfamiliar, counterfactually-trained agents should generalize better by leveraging learned notions of intervention quality. We validate this through a method we call Interruptible Collaborative Roleplayer (ICR), where we withhold common ground-based rewards during training and show that such agents still achieve greater convergence than sophisticated LLM-agent training baselines, suggesting they have internalized collaboration principles transferable across partners and task to “in-the-wild” settings. Our work advances the state of the art in LLM-based collaborative agents through the following contributions [^1] Our code is available at https://github.com/csu-signal/ICR:
@@
  • Theoretical insights demonstrating why standard reinforcement learning and preference alignment algorithms (e.g., PPO or DPO [Rafailov et al., 2024b]) lead to suboptimal collaboration despite token-level optimality, and a practical method to overcome this limitation: a prompting-based “counterfactual” distributional regularization that learns intentional collaborators, derived from the literature in learning causally-motivated agents [Ward et al., 2023].
@@ 2 Related Work
While interruptibility has been studied in safety-critical RL [Orseau and Armstrong, 2016, Hadfield-Menell et al., 2017], it is equally vital in collaborative dialogue, where agents must discern whether interventions aid or hinder shared understanding [Grice, 1975, Sutton and Rao, 2024]. Prior work has explored these ideas in adversarial or game-theoretic contexts [Langlois and Everitt, 2021, Ward et al., 2023], but less so in multi-party deliberative language settings [Nath et al., 2025a, Obiso et al., 2025]. We extend this by training collaborator agents that are counterfactually robust; they update their beliefs when interventions are helpful, while resisting misleading or misaligned input.
@@ 3 The Collaborator’s Dilemma
An AI collaborator that merely mimics behavior patterns or reflexively adopts suggestions may initially appear cooperative but will demonstrate poor robustness when faced with interventions that are noisy, irrelevant, or potentially misleading [Jaques et al., 2019]. Rather, it needs to develop what Ward et al. [2023] terms “intentionality,” or the capacity to autonomously evaluate interventions based on their causal impact on task outcomes rather than superficial plausibility. To address this limitation, we need a learning paradigm that enables collaborators to be partner-aware, or capable of adapting to specific intervention agents through selective incorporation of helpful suggestions while maintaining invariance to misleading ones—thereby developing the "intentionality" necessary for robust collaborative reasoning. Such a collaborator would maintain reasoned agency in the face of various intervention qualities, leading to more robust collaboration and better common ground convergence across diverse interaction scenarios. In other words, effective collaborators must remain safely interruptible [Orseau and Armstrong, 2016]. This is a delicate balance between receptive and robust that renders them open to incorporating valuable insights that genuinely contribute to task success, yet capable of maintaining their reasoning integrity when faced with misleading suggestions. This motivates our Interruptible Collaborative Roleplayer (ICR) learning algorithm.
@@ 4 Method: Interruptible Collaborative Roleplayer
Our counterfactual invariance approach directly addresses the suboptimality gap identified in Theorem 3.2. Initially during training of a collaborator policy $\pi_{C}^{CI}$ with counterfactual invariance regularization, the counterfactual KL divergence $\Delta_{\text{CF}}(\pi_{C}^{CI})=\mathbb{E}_{s,a^{I}}[D_{KL}(\pi_{C}^{CI}(\cdot|s,a^{I})\,\|\,\pi_{C}^{CI}(\cdot|s^{\text{CF}},a^{I}))]$ will be high as the policy has not yet learned to distinguish intervention quality, but decreases as training progresses and the policy acquires counterfactual robustness. As established in Theorem B.4, this directly bounds the suboptimality gap: $J_{\mathcal{M}}(\pi_{C}^{*})-J_{\mathcal{M}}(\pi_{C}^{CI})\leq\frac{2\gamma R_{max}}{(1-\gamma)^{2}}(\epsilon_{task}+C\cdot\Delta_{\text{CF}}(\pi_{C}^{CI}))$. Theoretically, as $\lambda_{\text{Intent}}\rightarrow\infty$, $\Delta_{\text{CF}}(\pi_{C}^{CI})$ approaches zero, making our policy’s performance approach that of the optimal policy $\pi_{C}^{*}$ (subject to task optimization constraints $\epsilon_{task}$). This theoretical guarantee connects directly to Lemma˜B.2 showing that while preference-aligned policies achieve Bellman optimality at both token and intervention levels in the underlying MDP, they remain suboptimal in the MAMDP due to failing to account for intervention quality. Our counterfactual invariance objective $\mathcal{J^{*}}(\theta_{C})$ bridges this gap by explicitly teaching collaborator LLM agents to distinguish between interventions based on their causal impact on task outcomes during training, rather than merely using their in-context learning capacity. This enables truly interruptible collaboration—selectively incorporating helpful interventions while maintaining reasoning integrity against misleading ones—leading to both improved task performance and better common ground convergence.
@@ 5 Experimental Design
For training data, we first collect MAMDP interaction trajectories (as defined in Sec. 3) on these two domains over 15 turns [^6] This roughly reflects the true distribution of back and forth interactions in the original DeliData task. using a high-capacity LLM (GPT-4o [OpenAI et al., 2024]) to roleplay both the intervener and the collaborator agents in each task. See Figs. 2 and 5 for prompts. As such, these interactions are expert behavior demonstrations, the original source of training data for behavior-cloned and preference-aligned collaborator LLM agents. For evaluation, all trained ICR collaborators and competing baselines are first deployed following the MAMDP interaction in the expert data collection, and then evaluated, primarily on their ability to reach consensus during the collaboration. In all cases, we use a fixed intervention agent—an instance of GPT-4o prompted with the same system prompt in all evaluation runs—with $T=0$ and top-$p$ of 0.9 for sampling. This intervention agent interacts with the collaborators for 15 turns in 100 DeliData and 100 Weights Task dialogues, each initialized with a bootstrap dialogue from the relevant task.
@@
The “no-press” setting explores whether collaborator agents can achieve objective alignment on accurate decisions without explicit modeling of how interventions influence common ground formation. Therefore, to control for language interpretability, rather than using full natural language, collaborator agents act over a discrete space of structured beliefs—allowing us to evaluate grounded reasoning without requiring fluency. In the Weights Task, collaborators express beliefs as symbolic propositions over block weights (i.e., green > red, blue = 10g), while in DeliData, agents select from predefined stances toward questions of which cards to flip: support, oppose, unsure, or consider_later. Agents are trained independently using only task-specific proxy rewards: factual accuracy in the Weights domain, and logically aligned card-checking in DeliData (e.g., +1 for supporting parts of a correct solution—a vowel or odd number, -1 for incorrect support, +0.5 for justified uncertainty). Using a proxy reward during training is intuitive as well as fair for baseline comparisons, since otherwise RL-based agent training is prone to reward hacking [^7] In fact, in our preliminary experimentation we found that rewarding agents with a consensus signal is counterproductive and often leads to reduced task-specific utility or correctness over propositions. [Strathern, 1997, Amodei et al., 2016]. In the no-press condition, evaluation follows an exact reward function $R(a^{c})$ that can be directly computed from the discrete solutions chosen by collaborators. For the full-press evaluation, we use an LLM-Judge [Zheng et al., 2023, Lambert et al., 2025]-based reward $R(s,a^{c})$, where $s$ is some dialogue context with the intervention present and $a^{c}$ are discrete actions inferred by the LLM-Judge based on the collaborator utterances in context.
@@ 6 Results and Analysis
We report “full-press” and “no-press” results for both tasks. For Weights Task (WTD), we report accuracy scores (ACC), a composite metric that multiplies the percentage of correct propositions by the total size of the common ground, rewarding both factual accuracy and common-ground convergence, while penalizing trivial solutions. For example, ACC of 14 indicates that the collaborator agents were able to recover 14 out of a 37 total theoretically possible propositions at the end of the collaboration, adjusted for correctness. For DeliData, we report both accuracy (ACC)—a task-specific fine-grained score [Karadzhov et al., 2024] based on the final submission (after $N=15$ turns)—and common ground gain (CG), defined as the net increase in unique solution types introduced during the dialogue beyond those initially proposed. Specifically, we subtract the number of unique solution frameworks (e.g., $Odd$, VowelVowel, $Odd+Vowel$ in DeliData) initially proposed by the collaborator agents from the total number of distinct solutions considered throughout the entire dialogue. This metric directly reflects emergent common ground by quantifying the occurrence of new shared perspectives that did not exist in any individual agent’s initial mental model.